OpenStudy (anonymous):

What is the value of the expression –4x3 when x = –2?

OpenStudy (***[isuru]***):

is it \[(-4) \times 3x\] or \[(-4)\times x ^{3}\]

OpenStudy (anonymous):

idk..

OpenStudy (anonymous):

Look : 

OpenStudy (anonymous):

So have : 4 * (-2) = ? ?^ 3 ( I said 2) in up !

OpenStudy (anonymous):

@xXPhobiax_X : Got it ?:)

OpenStudy (anonymous):

no..O.O

OpenStudy (anonymous):

k ! First answer to 4 * (-2) = ?

OpenStudy (anonymous):

IDK

OpenStudy (anonymous):

4 * 2 = ?

OpenStudy (anonymous):

wat the hell is *

OpenStudy (anonymous):

K ! Means \[\times \]

OpenStudy (anonymous):

well you could have just done that instead of getting all technical with *

OpenStudy (anonymous):

Yes ! Now give another to * .! And can you answer to : 4 \[\times(-2)\]

OpenStudy (anonymous):

ok im seriously confused can just tell me the answer :|

OpenStudy (anonymous):

OK ! Look we have : x=-2 so : -4x^3(^means power !) So have : (-4) * (-2) = 8 and 8 ^ 3 = 512

OpenStudy (anonymous):

thats not one of the selections

OpenStudy (anonymous):

What is the value of the expression –4x3 when x = –2? A. –24 B. –32 C. 24 D. 32

OpenStudy (anonymous):

k ! Its C .

OpenStudy (***[isuru]***):

if it was\[-4x ^{3}\] it says \[(-4) \times x^{3}\] not \[[(-4)\times x ]^{3}\]

OpenStudy (anonymous):

I thought 3 is power !:)

OpenStudy (***[isuru]***):

yep! ur answer sheets proves my expression

OpenStudy (***[isuru]***):

First of all the expression is \[(-4) \times x ^{3}\]

OpenStudy (***[isuru]***):

and they r asking the value of this expression when x is -2 which is \[(-4) \times x ^{3}\] = \[(-4) \times (-2)^{3}\] = (-4) x (-8) = 32 did u get it .... ?